OpenStudy (anonymous):

find the slope of a line perpendicular to the line containing points (-9,2) and (-8,-9)

OpenStudy (igreen):

Okay, first let's find the slope of the line that contains those points. Plug (-9, 2) and (-8, -9) into the slope formula. \(m = \dfrac{y_2-y_1}{x_2-x_1}\) \(m = \dfrac{-9-2}{-8+9}\) Simplify